OpenStudy (anonymous):

I need help finding the exact value of cos^2pi/8-sin^2pi/8

OpenStudy (badhi):

use the identity, $$\cos (2x)=\cos^2(x)-\sin^2(x)$$

OpenStudy (anonymous):

Simplify the following: (cos(pi))^2/8-sin(pi)^2/8 Evaluate cos(pi)^2. cos(pi)^2 = 1: 1/8-sin(pi)^2/8 Evaluate sin(pi)^2. sin(pi)^2 = 0: 1/8-0/8 Any number times zero is zero. 0/8 = 0: 1/8 Look for the additive identity. 1/8+0 = 1/8: Answer: 1/8

OpenStudy (anonymous):

Sorry I wrote the problem wrong its actually cos^2(pi/8)-sin^2(pi/8)

OpenStudy (badhi):

Still you can use the above identity to find the answer

OpenStudy (anonymous):

Can you give me the first step please I'm still lost

OpenStudy (anonymous):

okaii so these are the steps for the problem: cos^2pi/8-sin^2pi/8 you can you the same methods and steps(similar) to do the other one you want to do? any questions or doubts you still have please post them :)

OpenStudy (badhi):

@some_someone jinxed401 is mentioning about $$\cos^2(\pi/8)-\sin^2(\pi/8)$$ not $$\frac{\cos^2(\pi)}{8}-\frac{\sin^2(\pi)}{8}$$ @jinxed401 As i said, consider the identity, $$\cos(2x)=\cos^2(x)-\sin^2(x)$$ if you substitute x=pi/8 what would the above equation be?

OpenStudy (anonymous):

\[\cos(2\times \pi/8 ) =\cos^2(\pi/8)-\sin^2(\pi/8)\] I"m still not sure where to go here

OpenStudy (raden):

the finally it can be cos(2pi/8) = cos(pi/4) = cos45 degrees = ... ?

OpenStudy (raden):

45 degrees was the special angle in trigono.. u have to memorize the values of sin, cos, tan for the special angles

OpenStudy (anonymous):

cos45 is \[(\sqrt{2}/2)\]

OpenStudy (raden):

yes, that;s right

OpenStudy (anonymous):

but I'm not seeing where you got 45 degrees from though because (pi/8) is does not come out to be (pi/4)

OpenStudy (raden):

2/8 = 1/4 and 1 pi radian = 180 degrees so, 1/4 pi = 1/4 * 180 degrees = 45 degrees

OpenStudy (raden):

it just converting unit radian to degree

OpenStudy (anonymous):

Yes but where are you getting (1pi/8)

OpenStudy (raden):

hoho... ur original problem is cos^2pi/8-sin^2pi/8 so, we change this by identity what our friend said above ^ thus, cos^2pi/8-sin^2pi/8 = cos(2pi/4) just doing for right side, dont look for left side because it has passed away :)

OpenStudy (anonymous):

you are an amazing person

OpenStudy (raden):

i just continue it :)